# In-Context Multi-Operator Learning with DeepOSets

### v1

@@ 4 Universality of DeepOSets for In-Context Operator Learning
Given $\varepsilon>0$, $C>1$ and a compact set $\mathcal{K}\subset\mathcal{O}$, there exist integers $N,M\geq 1$, $\delta>0$, fixed discretizations $\textbf{x}=\{x_{1},...,x_{k}\}$ of $K_{1}$ and $\{y_{1},...,y_{n}\}$ of $K_{2}$, and continuous functions $\Psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{M}$, $\rho:\mathbb{R}^{k+N}\rightarrow\mathbb{R}^{M}$, and $\Phi:\mathbb{R}^{k+n}\rightarrow\mathbb{R}^{N}$, such that
$\sup_{G\in\mathcal{K}}\>\left|\,G(u_{q})(y)-\Psi(y)\cdot\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\right|<\varepsilon.$
for any $(\delta,C)$-discretization $\textbf{u}:=(u_{1},\cdots,u_{m})$ of $V$, any $u_{q}\in V$, and any $y\in K_{2}$.
@@
See the Appendix.∎
The continuous functions guaranteed by Theorem 4.2 can be approximated using neural networks by invoking standard universal approximation theorems for MLPs cybenko1989approximation; hornik1991approximation; pinkus1999approximation. Here it is important that we are averaging the outputs of $\Phi$ to ensure that the function $\rho$ only needs to be approximated on a compact set.
@@ 5 Experimental Results
Here, we demonstrate that the proposed DeepOSets architecture can learn the solution operator for multiple PDEs, with the correct PDE being disambiguated with the prompt examples, without any further weight updates. We consider two one-dimensional parametric PDEs and forward and inverse problems for each.
The Poisson boundary-value PDE reads:
$\frac{d^{2}u(x)}{dx^{2}}$ $=\,f(x)\,,\quad 0<x<1\,$, (2)
with $u(0)=u_{0}$ and $u(1)=u_{1}$. The forward problem is to find the solution $u(x)$ given the forcing function $f(x)$. The inverse or control problem is to find the forcing function $f(x)$ that produces the given trajectory $u(x)$. Very importantly, the boundary values $u_{0}$ and $u_{1}$ are unknown and must be learned in-context.
The reaction-diffusion boundary-value PDE is:
$a\frac{d^{2}u(x)}{dx^{2}}+k(x)u(x)=c\,,\quad 0<x<1\,$, (3)
with $u(0)=u_{0}$ and $u(1)=u_{1}$. The forward problem is to find the solution $u(x)$ given the coefficient function $k(x)$. The inverse problem is to find the coefficient function $k(x)$ that produces the given solution $u(x)$. Here, the coefficients $a$ and $c$ and the boundary values $u_{0}$ and $u_{1}$ are all unknown and must be learned in-context.
(a)         Forward Poisson
    $\frac{d^{2}{\color[rgb]{1,0,0}G(u)(x)}}{dx^{2}}={\color[rgb]{0,0,1}u(x)}$, given $u(0),u(1)$
(b)         Inverse Poisson
    $\frac{d^{2}{\color[rgb]{0,0,1}u(x)}}{dx^{2}}={\color[rgb]{1,0,0}G(u)(x)}$
, given $u(0),u(1)$
(c)      Forward Reaction-Diffusion
   $a\frac{d^{2}}{dx^{2}}{\color[rgb]{1,0,0}u(x)}+{\color[rgb]{0,0,1}k(x)}{\color[rgb]{1,0,0}u(x)}=c$, given $a$, $c$
(d)      Inverse Reaction-Diffusion
   $a\frac{d^{2}}{dx^{2}}{\color[rgb]{0,0,1}u(x)}+{\color[rgb]{1,0,0}k(x)}{\color[rgb]{0,0,1}u(x)}=c$, given $a$, $c$
Figure 2: In-context PDE solution operator learning with the DeepOSets architecture. In each case, the top plot contains the parameter functions while the bottom plot contains the corresponding solution functions. The gray lines indicate in each case the parameter-solution function pairs given in the prompt. The blue line indicates the query parameter function, while the red line indicates the corresponding solution predicted by the trained DeepOSets model. The black line indicates the exact solution corresponding to the query, using the appropriate PDE. The DeepOsets model is trained to learn the four PDE problems simultaneously, with coefficients sampled randomly.
@@
All PDE coefficients and boundary values are chosen randomly, so that the number of solution operators involved is in fact infinite. The physical domain in all cases is the interval $[0,1]$, which is discretized uniformly into 100 points. The prompts during both training and testing contain 4 example pairs of parameter and solution functions. The MLP $\phi$ in the DeepOSets contains 2 hidden layers of 400 neurons each, and the latent vector encoding the prompt data has size 500. The pooling operation is the sample mean. The branch and trunk networks in the DeepOSets contain hidden layers of 200, 400, and 200 neurons with intermediate layer of 150 neurons. All neural network use ReLU as the activation function.
@@
To generate the training data, for each of the four problems (Poisson/reaction-fiffusion, forward/inverse), 100 parameter functions are sampled independently from a Gaussian process with the squared-exponential kernel (discretized on the uniform 100-point grid) with variance 2 and length scale $0.5$. For each parameter function and PDE, the boundary values $u_{0}$ and $u_{1}$ are sampled uniformly in the interval [1,1][-1,1]. In addition, in the case of the reaction-diffusion PDE, the coefficients $a$ and $c$ are sampled uniformly in the intervals $[0.025,0.075]$ and [2,2][-2,2], respectively. Then the PDE is solved for the given parameter function, boundary values, and coefficients to obtain the groundtruth solutions. The model was trained by minimizing the mean squared error loss between predicted and groundtruth solutions over a mini-batch of 100 operators sampled randomly from the training data at each iteration, using the Adam optimizer with learning rate 1e-3, for 50k iterations. Training took 14min, achieving a training loss 8.4e-3.
Sample results are shown in Figure 2. At inference time, DeepOSets can perform accurate in-context learning, without knowing which PDE, inverse or forward problem, or the value of the equation coefficients and boundary conditions are. All these unknowns are correctly disambiguated from the data in the prompt, and the solution corresponding to the query parameter is accurately predicted, without any weight updates.
@@ 6 Conclusion
We have presented DeepOSets for in-context multi-operator learning, a novel non-autoregressive, non-attention-based approach to learning solution operators of multiple parametric PDEs. By combining DeepSets for set learning with DeepONets for operator learning, DeepOSets addresses key limitations of existing approaches while maintaining strong in-context learning capabilities. The proposed architecture offers several advantages: fully parallel processing of prompts enables faster training and inference compared to sequential methods; linear complexity in prompt length provides better scalability than quadratic self-attention mechanisms; and built-in permutation-invariance provides beneficial inductive bias for training efficiency. We prove that DeepOSets are universal uniform approximators over a class of operators, which we believe is the first result of its kind to be published in the Scientific Machine Learning area. Our experiments on Poisson and reaction-diffusion problems demonstrate that DeepOSets successfully disambiguates between multiple PDEs, forward and inverse problems, and unknown equation coefficients and boundary conditions purely from prompted examples. The model achieves accurate predictions without weight updates at inference time, requiring only 14 minutes of training on all four problems simultaneously. Future work will extend DeepOSets to higher-dimensional PDEs, more diverse equation families, and longer prompts. DeepOSets establishes a computationally efficient foundation for building general-purpose PDE solvers capable of in-context adaptation, contributing toward foundation models for scientific computing.
Appendix
Let us begin by proving Lemma 4.1 showing that $(\delta,C)$-discretizations always exist.
@@
Let $K\subset X$ be compact, and let $\delta>0$ be given. Choose a maximal $\delta$-separated subset of $K$, i.e., a subset $\mathcal{X}_{\delta}$ such that $d(x_{i},x_{j})\geq\delta$ for all $i\neq j$, and which is maximal with respect to this property. It is clear that the balls of radius $\delta$ around the points $x_{i}$ give a minimal covering of $K$, and so by compactness $\mathcal{X}_{\delta}$ must be a finite set. We claim that $\mathcal{X}_{\delta}$ is a $(\delta,C)$-covering for any $C>1$.
First, note that any ball of radius $\delta/2$ contains at most one point of $\mathcal{X}_{\delta}$. Indeed, fix $z\in K$ and suppose $x_{i},x_{j}\in\mathcal{X}_{\delta}$ both lie in $B(z,\delta/2)$ with $i\neq j$. Then, by the triangle inequality,
$d(x_{i},x_{j})\;\leq\;d(x_{i},z)+d(z,x_{j})\;<\;\frac{\delta}{2}+\frac{\delta}{2}\;=\;\delta,$
which contradicts the $\delta$–separation of $\mathcal{X}_{\delta}$. Hence, for every $z\in K$,
$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(z,\delta/2)\}\bigr|\leq 1.$
Next, since balls of radius $\delta$ around the points in $\mathcal{X}_{\delta}$ cover $K$, for every $y\in K$ there exists some $x_{j}\in\mathcal{X}_{\delta}$ with $d(y,x_{j})<\delta$, i.e. $x_{j}\in B(y,\delta)$. Thus, for all $y\in K$,
$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(y,\delta)\}\bigr|\geq 1.$
Combining these two observations, we obtain, for all $x,y\in K$,
$\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(x,\delta/2)\}\bigr|\;<\;C\,\bigl|\{x_{i}\in\mathcal{X}_{\delta}:x_{i}\in B(y,\delta)\}\bigr|,$
for any $C>1$. Hence $\mathcal{X}_{\delta}$ is a $(\delta,C)$–discretization of $K$. ∎
Next, we state a Lemma which we will use in the proof of Theorem 4.2.
Lemma 6.1.
@@
(of Lemma 6.1) Since $V$ is compact, by the Arzela-Ascoli theorem it is also equicontinuous, and so there exists $\delta>0$ such that if $|x-y|<\delta$, then $|f(x)-f(y)|<\varepsilon$ for all $f\in V$. Since $K$ is also compact, we can pick $y_{1},\cdots,y_{n}\in K$ such that $\{B(y_{i},\delta)\}_{i=1}^{n}$ form a covering of $K$. Finally, let $\{\varphi_{i}\}_{i=1}^{n}$ be a partition of unity subordinate to the balls $\{B(y_{i},\delta)\}_{i=1}^{n}$.
@@
is compact. Since $K_{2}$ is also compact, we can apply Lemma 6.1 to get that there exists a grid y of $K_{2}$, and $\{\gamma_{i}\}_{i=1}^{n}\subset C(K_{2})$, such that
@@
Since $\mathcal{K}$ is compact, it is equi-continuous, so there exists a $\delta$ such that if $\|u-v\|_{C(K_{1})}<25\delta/6$, then $\|G(u)-G(v)\|_{C(K_{2})}<\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})$ for every $G\in\mathcal{K}$. By Lemma 6.1, there exists $\textbf{x}=\{x_{1},\cdots,x_{k}\}\in K_{1}$ such that
@@
We then have that for any $G\in\mathcal{K}$,
@@
$\|G(u_{q})-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K)}$
$\leq\|G(u_{q})-G(u_{q})(\textbf{y})\cdot\Psi\|_{C(K)}+\|G(u_{q})(\textbf{y})\cdot\Psi-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K)}$
@@
Finally, we need to show that $\rho$ and $\Phi$ can be chosen continuous. It is clear from the definition that $\Phi$ is continuous. To show that $\rho$ is also continuous, we need to show that the denominator

### v2

@@ 5 Universality of DeepOSets for In-Context Operator Learning
Given $\varepsilon>0$, $C>1$ and a compact set $\mathcal{K}\subset\mathcal{O}$, there exist integers $N,M\geq 1$, $\delta>0$, fixed discretizations $\textbf{x}=\{x_{1},...,x_{k}\}$ of $K_{1}$ and $\textbf{y}=\{y_{1},...,y_{n}\}$ of $K_{2}$, and continuous functions $\Psi:\mathbb{R}^{d}\!\rightarrow\mathbb{R}^{M}$, $\rho:\mathbb{R}^{k+N}\!\rightarrow\mathbb{R}^{M}$, $\Phi:\mathbb{R}^{k+n}\!\rightarrow\mathbb{R}^{N}$, such that
$\sup_{G\in\mathcal{K}}\>\Bigg|\,G(u_{q})(x_{q})~-~\left.\Psi(x_{q})\cdot\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\right|<\varepsilon\,$, (3)
for any $(\delta,C)$-discretization $\textbf{u}:=(u_{1},\cdots,u_{m})$ of $V$, any $u_{q}\in V$, and any $x_{q}\in K_{2}$.
@@
For the full detailed proof, please see the Appendix.
Here, we present an overview of the proof, in which the approximation error is decomposed into two inequalities. The compactness assumption enables this decomposition to hold uniformly over the class of continuous operators by reducing the problem to finite-dimensional approximations via discretizations on the grids $\mathbf{x}$ and $\mathbf{y}$.
For the first inequality, using a partition of unity Lemma (in Appendix), we show that there exists a grid y of $K_{2}$, and $\{\gamma_{i}\}_{i=1}^{n}\subset C(K_{2})$, such that
for every $G\in\mathcal{K}$ and $u_{q}\in V$.
For the second inequality, we show that there exists a grid x of $K_{1}$ such that
$G(u_{q})(\textbf{y}):=(G(u_{q})(y_{1}),\cdots,G(u_{q})(y_{n}))$
can be approximated well by functions of the form,
where $u_{1},...,u_{m}$ is a $(\delta,C)$-discretization of $V$.
We construct $\Phi$ using a partition-of-unity argument to be localized/subordinate over a discretization of the input space, and define $\rho$ by forming normalized local averages of the corresponding operator output samples evaluated on the output grid and recombining them into a continuous finite-dimensional approximation.
We then let $\Psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$, (here $n=M)$, be defined as,
$\Psi(x)=\{\gamma_{i}(x)\}_{i=1}^{n}\,.$
Taking the dot product with (4) and applying the triangle inequality together with the two bounds yields the desired approximation inequality.
Finally, we verify that the constructions $\Phi$ and $\rho$ are continuous. Continuity of $\Phi$ is immediate from its definition, while continuity of $\rho$ follows from uniform lower bounds on the normalization factors ensured by the balanced $(\delta,C)$-discretization of $V$. ∎
The continuous functions guaranteed by Theorem 5.1 can be approximated using neural networks by invoking standard universal approximation theorems for MLPs ((Cybenko, 1989; Hornik, 1991; Pinkus, 1999). Here it is important that we are averaging the outputs of $\Phi$ to ensure that the function $\rho$ only needs to be approximated on a compact set.
@@ 6 Experimental Results
# Problem Description Differential Equations Parameters Conditions QoIs
1 Forward IVP 1 $\frac{d}{dt}u(t)=a_{1}c(t)+a_{2}$ for $t\in[0,1]$ $a_{1},a_{2}$ $u(0),c(t),t\in[0,1]$ $u(t),t\in[0,1]$
2 Inverse IVP 1 $u(t),t\in[0,1]$ $c(t),t\in[0,1]$
3 Forward IVP 2 $\frac{d}{dt}u(t)=a_{1}c(t)u(t)+a_{2}$ for $t\in[0,1]$ $a_{1},a_{2}$ $u(0),c(t),t\in[0,1]$ $u(t),t\in[0,1]$
4 Inverse IVP 2 $u(t),t\in[0,1]$ $c(t),t\in[0,1]$
5 Forward IVP 3 $\frac{d}{dt}u(t)=a_{1}u(t)+a_{2}c(t)+a_{3}$ for $t\in[0,1]$ $a_{1},a_{2},a_{3}$ $u(0),c(t),t\in[0,1]$ $u(t),t\in[0,1]$
6 Inverse IVP 3 $u(t),t\in[0,1]$ $c(t),t\in[0,1]$
7 Forward Poisson BVP $\frac{d^{2}}{dx^{2}}u(x)=f(x)$ for $x\in[0,1]$ $u(0),u(1)$ $f(x),x\in[0,1]$ $u(x),x\in[0,1]$
8 Inverse Poisson BVP $u(x),x\in[0,1]$ $f(x),x\in[0,1]$
9 Forward linear reaction-diffusion BVP $a\frac{d^{2}}{dx^{2}}u(x)+k(x)u(x)=c$ for $x\in[0,1]$ $u(0),u(1),a,c$ $k(x),x\in[0,1]$ $u(x),x\in[0,1]$
10 Inverse linear reaction-diffusion BVP $u(x),x\in[0,1]$ $k(x),x\in[0,1]$
11 Forward nonlinear reaction-diffusion BVP $a\frac{d^{2}}{dx^{2}}u(x)+k(x)u(x)^{3}=c$ for $x\in[0,1]$ $u(0),u(1),a,c$ $k(x),x\in[0,1]$ $u(x),x\in[0,1]$
12 Inverse nonlinear reaction-diffusion BVP $u(x),x\in[0,1]$ $k(x),x\in[0,1]$
Table 1: Summary of tasks. Two benchmarks are considered: in the first, only the first six tasks, corresponding to the initial-value problems (IVP), are considered; in the second, all twelve tasks, consisting of both IVPs and boundary-value problems (BVP), are considered jointly.
We demonstrate that the DeepOSets architecture successfully learns the solution operator for various initial-value problems (IVP) and boundary-value problems (BVP). Our evaluation spans 12 distinct tasks, which are summarized in Table 1. This table and the terminology “condition” and “quantity of interest (QoI)” are adapted from Yang et al. ((2023b). There are in fact an infinite number of solution operators in this experiment, corresponding to different settings of equation coefficients and boundary values. The prompt examples allow the model to disambiguate the specific operator in-context, without requiring additional weight updates. We fix $u(0)=0$ in all IVPs (tasks 1–6). In the BVPs (tasks 7–12), the boundary values $u(0)$ and $u(1)$ are chosen randomly and are unknown to the model. To evaluate how task range and diversity influence performance, we consider two benchmarks: in the first, the DeepOSets model is trained and tested only on the 6 boundary-value problems, while in the second, the model is trained and tested on all 12 initial and boundary value problems. As will be seen below, training on the 12 tasks indeed produces a more accurate multi-operator learner.
@@ 12 Inverse nonlinear reaction-diffusion BVP $u(x),x\in[0,1]$ $k(x),x\in[0,1]$
In all experiments, equation coefficients and boundary condition values are chosen randomly. The physical domain in all cases is the interval $[0,1]$, which is discretized uniformly into 100 points. The MLP $\phi$ in the DeepOSets contains 2 hidden layers of 500 neurons each, and the latent vector encoding the prompt data has size 500. The pooling operation is the sample mean. The branch and trunk networks in the DeepOSets model contain hidden layers of 200, 500, and 200 neurons. The pooled latent space is 2000 with mean pooling. All neural network use ReLU as the activation function.
@@
To generate the training data for each of the problems, 100 parameter functions are sampled independently from a Gaussian process with the squared-exponential kernel (discretized on the uniform 100-point grid) with variance 2 and length scale $0.5$. For each boundary-value problem, the boundary values $u(0)$ and $u(1)$ are sampled uniformly in the interval [1,1][-1,1]. For the reaction-diffusion problems, the coefficients $a$ and $c$ are sampled uniformly in the intervals $[0.025,0.075]$ and [2,2][-2,2], respectively. Then the equation is solved for the given parameters to obtain the ground truth solutions. The model was trained by minimizing the mean-squared error loss between predicted and ground truth solutions over a mini-batch of 100 operators sampled randomly from the training data at each iteration, using the Adam optimizer with learning rate 1e-3, for 50k iterations. Training the 6-task DeepOSets took only 12min, achieving a training loss 8.4e-3, while the training time of the 12-task DeepOSets was only 38 minutes, achieving a training loss of 4.8e-3, on an Nvidia 4060 Ti GPU with 16GB memory. We contrast this to the complex transformer-based architecture in Yang et al. ((2023b), which requires many hours of training on similar computing hardware.
Sample results are shown in Figure 2, while MSE values on test data from operators not seen during training are displayed in Table 2. DeepOSets achieves accurate predictions without weight updates at inference time, without knowing which task or the values of the equation coefficients and boundary conditions. All these unknowns are correctly disambiguated from the examples in the prompt. By comparing the 6-task and 12-task errors in Table 2, we observe the interesting fact that the errors are almost universally smaller with the 12-task DeepOSets. Considering all 12 tasks simultaneoulsy could be expected to be more challenging due to the larger number of equations but, on the other hand, it presents a larger diversity of solution behaviors. Our experiment shows that more diversity of solution operators in the training data is indeed helpful in training a better model that learns to disambiguate the multiple operators more effectively in-context.
(a) Forward Poisson BVP
(b) Inverse Poisson BVP
(c) Forward reaction-diffusion BVP
(d) Inverse reaction-diffusion BVP
(e) Forward nonlinear reaction-diffusion BVP
(f) Inverse nonlinear reaction-diffusion BVP
Figure 2: Sample results for boundary-value problem test data obtained by the 12-task model. In each case, the top plot contains the “condition” (input) functions while the bottom plot contains the corresponding “QoI” (output) functions. The gray lines indicate in each case the condition-QoI function pairs given in the prompt. The blue line indicates the query parameter function, while the red line indicates the corresponding output of the trained DeepOSets model. The black line indicates the groundtruth output corresponding to the query.
Test Problem 6-task DeepOSets 12-task DeepOSets
Forward IVP 1 2.65e-03 $\pm$ 9.30e-03
Inverse IVP 1 1.15e-02 $\pm$ 2.71e-02
Forward IVP 2 1.65e-03 $\pm$ 8.39e-03
Inverse IVP 2 1.44e-02 $\pm$ 3.86e-02
Forward IVP 3 1.19e-02 $\pm$ 6.99e-02
Inverse IVP 3 1.40e-02 $\pm$ 9.15e-02
Forward Poisson BVP 3.04e-03 $\pm$ 1.91e-02 4.87e-04 $\pm$ 1.09e-03
Inverse Poisson BVP 4.60e-03 $\pm$ 1.35e-02 3.84e-03 $\pm$ 1.20e-02
Forward linear reaction-diffusion BVP 2.11e-03 $\pm$ 4.25e-03 6.49e-04 $\pm$ 9.75e-04
Inverse linear reaction-diffusion BVP 1.45e-02 $\pm$ 6.01e-02 7.71e-03 $\pm$ 2.75e-02
Forward nonlinear reaction-diffusion BVP 2.77e-03 $\pm$ 5.69e-03 1.71e-03 $\pm$ 3.32e-03
Inverse nonlinear reaction-diffusion BVP 1.19e-02 $\pm$ 3.51e-02 1.33e-02 $\pm$ 5.42e-02
Table 2: Mean Squared Error (MSE) values achieved by the 6-task and 12-task models on test data from each of the 12 tasks individually.
Next, we evaluated the ability of the DeepOSets framework to extrapolate to sample sizes in inference time that were not seen at training time. Indeed, the promised advantages of the DeepOSets approach are its permutation invariance and flexibility regarding the number of examples in the prompt. To evaluate the extrapolation capability, we used different number of in-context examples ranging from $N=3$ to $N=20$ at inference time. As illustrated in Figure 3, the model successfully generalizes to context sizes significantly larger than those seen in training, and we observe a consistent decrease in MSE as the number of context examples increases. This behavior confirms that the model is not overfitting to the specific training sequence length; rather, it effectively performs in-context learning, leveraging additional data points to refine the operator approximation, demonstrating robust generalization.
(a)   Forward problems
(b)   Inverse problems
Figure 3: Zero-shot generalization to larger context sizes. The model was trained with a fixed context size of $N=4$. During inference, we evaluate performance on context sizes ranging from 3 to 20. The results show a clear trend where providing more examples decreases the MSE, indicating the model successfully leverages additional data.
@@ 7 Conclusion
We have presented DeepOSets for in-context multi-operator learning, a novel non-autoregressive, non-attention-based approach to learning solution operators of multiple parametric differential equations. By combining DeepSets for set learning with DeepONets for operator learning, DeepOSets addresses key limitations of existing approaches while maintaining strong in-context learning capabilities. The proposed architecture offers several advantages: fully parallel processing of prompts enables faster training and inference compared to sequential methods; linear complexity in prompt length provides better scalability than quadratic self-attention mechanisms; and built-in permutation-invariance provides beneficial inductive bias for training efficiency. We prove that DeepOSets are universal uniform approximators over a class of operators, which we believe is the first result of its kind to be published in the field of multi-operator learning. Our experiments on different tasks from initial-value and boundary-value problems demonstrate that DeepOSets successfully uses the prompted examples to disambiguate between multiple operators corresponding to forward and inverse problems and unknown equation coefficients and boundary conditions. The experiments also showed that using a more diverse training set with more solution operators was beneficial to accuracy of in-context learning with the DeepOSets model. DeepOSets establishes a computationally efficient foundation for building general-purpose differential equation solvers capable of in-context adaptation, contributing towards establishing foundation models for scientific computing.
Appendix A Proofs of Theorem 5.1 and Lemma 3.2
Let us begin by proving Lemma 3.2 showing that $(\delta,C)$-discretizations always exist for any compact metric space $V$.
@@
Let $V\subset X$ be compact, and let $\delta>0$ be given. Choose a maximal $\delta$-separated subset of $V$, i.e., a subset $\mathcal{X}_{\delta}$ such that $d(u_{i},u_{j})\geq\delta$ for all $i\neq j$, and which is maximal with respect to this property. It is clear that the balls of radius $\delta$ around the points $u_{i}$ give a minimal covering of $V$, and so by compactness $\mathcal{X}_{\delta}$ must be a finite set. We claim that $\mathcal{X}_{\delta}$ is a $(\delta,C)$-covering for any $C>1$.
First, note that any ball of radius $\delta/2$ contains at most one point of $\mathcal{X}_{\delta}$. Indeed, fix $z\in V$ and suppose $u_{i},u_{j}\in\mathcal{X}_{\delta}$ both lie in $B(z,\delta/2)$ with $i\neq j$. Then, by the triangle inequality,
$d(u_{i},u_{j})\;\leq\;d(u_{i},z)+d(z,u_{j})\;<\;\frac{\delta}{2}+\frac{\delta}{2}\;=\;\delta,$
which contradicts the $\delta$–separation of $\mathcal{X}_{\delta}$. Hence, for every $z\in V$,
$\bigl|\{u_{i}\in\mathcal{X}_{\delta}:u_{i}\in B(z,\delta/2)\}\bigr|\leq 1.$
Next, since balls of radius $\delta$ around the points in $\mathcal{X}_{\delta}$ cover $V$, for every $v\in V$ there exists some $u_{j}\in\mathcal{X}_{\delta}$ with $d(v,u_{j})<\delta$, i.e. $u_{j}\in B(v,\delta)$. Thus, for all $v\in V$,
$\bigl|\{u_{i}\in\mathcal{X}_{\delta}:u_{i}\in B(v,\delta)\}\bigr|\geq 1.$
Combining these two observations, we obtain, for all $u,v\in V$,
$\bigl|\{u_{i}\in\mathcal{X}_{\delta}:u_{i}\in B(u,\delta/2)\}\bigr|\;<\;C\,\bigl|\{u_{i}\in\mathcal{X}_{\delta}:u_{i}\in B(v,\delta)\}\bigr|,$
for any $C>1$. Hence $\mathcal{X}_{\delta}$ is a $(\delta,C)$–discretization of $V$. ∎
Next, we state a partition of unity approximation Lemma for a compact subset of $C(K)$, which we will use in the proof of Theorem 5.1.
Lemma A.1.
@@
(of Lemma A.1) Since $V$ is compact, by the Arzela-Ascoli theorem it is also equicontinuous, and so there exists $\delta>0$ such that if $|x-y|<\delta$, then $|f(x)-f(y)|<\varepsilon$ for all $f\in V$. Since $K$ is also compact, we can pick $y_{1},\cdots,y_{n}\in K$ such that $\{B(y_{i},\delta)\}_{i=1}^{n}$ form a covering of $K$. Finally, let $\{\varphi_{i}\}_{i=1}^{n}$ be a partition of unity subordinate to the balls $\{B(y_{i},\delta)\}_{i=1}^{n}$.
@@
is compact. Since $K_{2}$ is also compact, we can apply Lemma A.1 to get that there exists a grid y of $K_{2}$, and $\{\gamma_{i}\}_{i=1}^{n}\subset C(K_{2})$, such that
@@
Since $\mathcal{K}$ is compact, it is equi-continuous, so there exists a $\delta$ such that if $\|u-v\|_{C(K_{1})}<25\delta/6$, then $\|G(u)-G(v)\|_{C(K_{2})}<\varepsilon/(2\cdot\|\Psi\|_{C(K_{2})})$ for every $G\in\mathcal{K}$. By Lemma A.1, there exists $\textbf{x}=\{x_{1},\cdots,x_{k}\}\in K_{1}$ such that
@@
where $\mathbf{u}\in\mathbb{R}^{k}$ and $\mathbf{v}\in\mathbb{R}^{n}$. We then have that for any $G\in\mathcal{K}$,
@@
$\|G(u_{q})-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K_{2})}$
$\leq\|G(u_{q})-G(u_{q})(\textbf{y})\cdot\Psi\|_{C(K_{2})}+\|G(u_{q})(\textbf{y})\cdot\Psi-\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right)\cdot\Psi\|_{C(K_{2})}$
@@
Since $\|.\|_{C(K_{2})}:=\sup_{x\in K_{2}}\|.(x)\|$, this result then holds for any $x_{q}\in K_{2}$, as desired. This completes the main inequality.
However, we still need to show that $\rho$ and $\Phi$ can be chosen continuous. It is clear from the definition that $\Phi$ is continuous. To show that $\rho$ is also continuous, we need to show that the denominator